# References.bib

### v7

@@ 4.2 Recursive definition step 1: $\theta(n)$
In Subsections 4.2 and 4.3 (the rest of this page) we aim to combine the two instances of the doubly-recursive definition of the Roman factorial:
$\lfloor n\rceil!=\begin{cases}n\lfloor n-1\rceil!&,\enspace n\in\mathbb{Z}^{+}\\[3.68027pt] \cfrac{\lfloor n+1\rceil!}{n+1}&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,,\end{cases}$ (Eq. 2.2.3)
@@
0!=1!=1$\lfloor$ 0$\rceil$!$\lfloor$-1$\rceil$!=1\,.⌊ 0 ⌉ ! = ⌊ - 1 ⌉ ! = 1 . (Eq. 4.1.1)
@@
In this way, we can see that there are two occurrences of a function in each case. That function is $+1$ when $n\in\mathbb{Z}^{+}$ and $-1$ when $n\in\mathbb{Z}^{-}\setminus\{-1\}$ (meaning $n$ is a negative integer except $-1$).
@@
$\lfloor n\rceil!=\begin{cases}n^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{Z}^{+}\\[3.68027pt] (n+1)^{\theta(n)}\lfloor n-\theta(n)\rceil!&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,,\end{cases}$ (Eq. 4.2.2)
@@
0!=1!=1$\lfloor$ 0$\rceil$!$\lfloor$-1$\rceil$!=1\,.⌊ 0 ⌉ ! = ⌊ - 1 ⌉ ! = 1 . (Eq. 4.1.1)
@@ 4.3 Recursive definition step 2: $\xi^{\prime}(n)$
$\xi^{\prime}(n)=\begin{cases}0&,\enspace n\in\mathbb{Z}^{+}\\ 1&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,.\end{cases}$ (Eq. 3.5.1)
@@
0!=1!=1$\lfloor$ 0$\rceil$!$\lfloor$-1$\rceil$!=1\,.⌊ 0 ⌉ ! = ⌊ - 1 ⌉ ! = 1 . (Eq. 4.1.1)
@@ 4.4 Roman-like definition step 1: $\eta(n)$
In Subsections 4.4, 4.5, 4.6 and 4.7 we aim to unite the two cases of the original definition of the Roman factorial, expressed as
@@ 4.5 Roman-like definition step 2: $\theta(n)$
For the second step, let’s observe that the factorial in the second case is inverted, while the factorial in the first case remains unchanged. In other words, the factorial of $n$ or $(-n-1)$ is raised to the power of 1 or $-1$, respectively:
@@
To express this change, we need a function for the exponent that equals 1 when $n$ is non-negative and -1 when $n$ is negative. Another way to describe this requirement is the output pattern $[-1,1,1]$.
@@
In the second case of Eq. 4.5.3, there is a negative sign before the $n$ in the term $(-n-1)$. This ensures that the factorial calculated is always positive, because if $n<0$, then $-n>0$. In contrast, in the first case, the $n$ inside the factorial is written without any sign (implying the positive sign).
@@
We are nearly at the point where both cases have identical expressions. The only remaining difference is the $-1$ inside the factorial of the second case, which we’ll address next.
@@ 4.7 Roman-like definition step 4: $\xi^{\prime}(n)$
Let’s investigate the $-1$ that appears only when $n\in\mathbb{Z}^{-}$. We can begin by expanding the first case as shown below:
@@
This marks the completion of one of our objectives. To sum up, we established a collection of foundational functions, which enabled us to express the definition of the Roman factorial concisely and efficiently. In the following section, we will introduce another set of F.F. which will be used alongside those defined in Section 3 in order to build advanced factorial definitions. For example, in Section 6, we aim to create a factorial similar to the Roman factorial, but expressed as a non-recursive product.
@@ 5.3 The function $Q(n)$
The function $Q(n)$ represents the $sign$ function, which exhibits the output pattern $[-1,0,1]$. It returns the sign of a number, like $\theta(n)$, but it equals 0 if that number is 0. Its piece-wise definition is this:
[…]
@@
Next, we will attempt to introduce another F.F. in the definition of $Q(n)$. Let’s start by examining the second case: we know that $10=\theta(0)-$ 1, so we can substitute that expression in the definition. Additionally, let’s rewrite the first case as $\theta(n)-0$, in order to highlight a subtraction from $\theta(n)$:
@@
This function is useful because it behaves identically to $sign(n)$, producing $+1$ or -$1$ for positive or negative inputs respectively, and 0 for the input 0. However, $Q(n)$ is defined using foundational functions, and henceforth, it will be used in place of $sign(n)$.
@@ 5.4 The function $Q^{\prime}(n)$
An easy approach to achieve this goal using the F.F. is to take the absolute value of $Q(n)$. Since $Q(n)$ has the output pattern $[-1,0,1]$, we see that $|Q(n)|$ will have the pattern $[1,0,1]$. That is our desired set of outputs, so we can define $Q^{\prime}(n)$ as the absolute value of $Q(n)$:
@@
Another approach is to subtract $\Theta(n)$ from 1, which is equally valid. Generally, the Boolean-like foundational functions presented in Section 5 can be combined in many ways to achieve a certain output pattern. Below you will find more alternative definitions of this function:
@@ 5.5 The function $\Psi(n)$
Additionally, we can interpret $\Psi(n)$ as a function that maps 00 to $1$ while leaving all other values of $n$ unchanged. The simplicity of this function makes it useful for various mathematical operations, but it’s most useful in defining the next F.F., $\Phi(n)$.
@@ 5.6 The function $\Phi(n)$
This works well, except for the case when $n=0$: $\Phi(0)=0^{0}$, which is problematic because $0^{0}$ is undefined [^16] The indeterminate form $0^{0}$ does not have a universally agreed-upon value, although it is given the value 1 in a lot of contexts [0tothe0]. More about $0^{0}$ in Addendum 9.6.. Fortunately we can solve this problem by using $\Psi(n)$. We end up with this universal definition of $\Phi(n)$:
@@ 5.7 Synopsis
In short, this section is about presenting another set of 5 foundational functions. They were all expressed in terms of other F.F., and along with those in Section 3 they constitute a set of Boolean-like functions that are instrumental to this paper.
In the following table, we list every foundational function defined in Sections 3 and 5 for $n\in\mathbb{R}$:
$\delta(n)$ $=\lfloor n\rfloor+0.5$ [,+,+[-,+,+][ - , + , + ] (Eq. 3.2.3)
@@
For example, the output pattern of $\xi^{\prime}(n)$ is $[1,0,0]$. That can be read as the binary number $100_{2}$, which is equal to 4 in base 10. Many F.F. exhibit this behavior, but those with outputs like $-1$ or $n$ are excluded.
@@
Table Tbl. 5.7.2: F.F. outputs in binary
@@ 6.1 Introduction
This section will focus on generalizing the $\prod$-product definitions of the Roman factorial, in the same manner as it was done in Section 4.
In Subsection 2.1, we introduced two formulations of the factorial function: one as a rising product and the other as a falling product. These expressions are reiterated below:
$n!=\prod_{k=1}^{n}k\,=1\cdot 2\cdot 3\cdots(n-2)(n-1)n\,,\quad n\in\mathbb{Z}^{+},$ (Eq. 2.1.1)
@@
We will follow a methodology akin to that presented in Section 4: utilizing F.F., we will formulate a piece-wise definition of the factorial. However, at present, we lack a $\prod$-product representation for negative Roman factorials, either as an ascending or descending product.
@@ 6.2 Rising product expression
$n$7 -6 -5 -4 -3 -2 -1
$\lfloor n\rceil!$ $\nicefrac{{1}}{{720}}$$\nicefrac{{1}}{{120}}$ $\nicefrac{{1}}{{24}}$$\nicefrac{{1}}{{6}}$ $\nicefrac{{1}}{{2}}$1 1
@@
Indeed, negative Roman factorials seem to be the product of consecutive reciprocals of integers, albeit with an offset. Despite the fractions getting smaller with each increment, since the terms start at $-1$ and end at $n$, we designate this a rising product. Disregarding the offset for now, this pattern can be expressed as follows:
$\lfloor n\rceil!=\prod_{k=1}^{-n}\cfrac{1}{-k}\enspace,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.2.1)
@@
There are two approaches to address this issue: one is to adjust the upper limit of the product, subtracting 1 from it to multiply by one less term. The other involves multiplying the existing product by $n$, which cancels out the term causing the offset. Both methods work equally well, although the first idea utilizes the empty product [^17] For more information about the empty product, check Addendum 9.2. when $n=-1$.
@@
$\lfloor n\rceil!=n\cdot\prod_{k=1}^{-n}\cfrac{1}{-k}\;,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.2.2)
@@ 6.3 Falling product expression
From our previous analysis, we observed that the negative Roman factorial of $n$ can be represented by a product consisting of $(n-1)$ terms. Thus, the upper and lower limits of the product must be set accordingly. A possible definition expressing the aforementioned products is as follows:
$\prod_{k=1}^{-n-1}\cfrac{1}{n+k}=\cfrac{1}{n+1}\cdot\cfrac{1}{n+2}\cdots\cfrac{1}{-2}\cdot\cfrac{1}{-1}\;,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.3.1)
While accurate and applicable to all negative integers, for $n=-1$, the use of the empty product is necessary again as the upper limit becomes smaller than the lower one. It is preferable to have a definition with the same lower product limit $k=0$ as in Eq. 2.1.2. So, we can multiply by one extra term and subsequently cancel it out, as demonstrated below:
$n\cdot\prod_{k=0}^{-n-1}\cfrac{1}{n+k}=\cfrac{n}{n}\cdot\cfrac{1}{n+1}\cdots\cfrac{1}{-2}\cdot\cfrac{1}{-1}\;,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.3.2)
Eq. 6.3.2 produces the desired values while also avoiding using the empty product when $n=-1$. Therefore, we will adopt this definition for the falling product expression:
@@
In summary, in Subsections 6.2 and 6.3, we’ve established two expressions describing negative Roman factorials as rising or falling products. They are consolidated below:
@@
$n\in\mathbb{Z}^{+}$ $\,\displaystyle\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Table Tbl. 6.4.1: Roman factorial as a rising or falling product (not including $n=0$)
@@
$n\in\mathbb{Z}^{+}_{0}$ $\,\displaystyle\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Table Tbl. 6.4.2: Roman factorial as a rising or falling product
Next, we notice the upper limits of the products. For $n\in\mathbb{Z}^{+}_{0}$, they are either $n$ or $n-1$, and for $n\in\mathbb{Z}^{-}$ they are $-n$ or $-n-1$.
Essentially, when $n$ becomes negative, a negative sign is applied to ensure that the upper limits remain positive numbers. As seen in Subsection 4.6, we can replace occurrences of $n$ and $-n$ with $|n|$. Consequently, the table is modified as follows:
@@
$n\in\mathbb{Z}^{+}_{0}$ $\,\displaystyle\prod_{k=1}^{|n|}k$ $\enspace\displaystyle\prod_{k=0}^{|n|-1}(n-k)$
@@
Table Tbl. 6.4.3: Roman factorial as a rising or falling product (generalization: step 1)
Unlike Section 4, this section generalizes two definitions simultaneously. While it’s possible to separate the procedure for the rising product from the falling one, the steps are identical: combining the two Roman factorial product definitions into a single table allows us to present the generalization process concurrently, unifying the expressions top to bottom.
@@ 6.5 $\prod$-product definition step 2: $\Phi(n)$
In this step, we will introduce the function $\Phi(n)$. As defined in Subsection 5.6, it takes the form:
@@
Table Tbl. 6.5.1: Roman factorial as a rising or falling product (generalization: step 2)

### v8

@@ 4.2 Recursive definition step 1: $\theta(n)$
In Subsections˜4.2 and 4.3 (the rest of this page) we aim to combine the two instances of the doubly-recursive definition of the Roman factorial:
[…]
@@
$\lfloor 0\rceil!=\lfloor-1\rceil!=1\,.$ (Eq. 4.1.1)
@@
In this way, we can see that there are two occurrences of a function in each case. That function is +1+1 when $n\in\mathbb{Z}^{+}$ and 1-1 when $n\in\mathbb{Z}^{-}\setminus\{-1\}$ (meaning $n$ is a negative integer except 1-1).
@@
[…]
@@
$\lfloor 0\rceil!=\lfloor-1\rceil!=1\,.$ (Eq. 4.1.1)
@@ 4.3 Recursive definition step 2: $\xi^{\prime}(n)$
$\xi^{\prime}(n)=\cfrac{1-\theta(n)}{2}=\begin{cases}0&,\enspace n\in\mathbb{Z}^{+}\\ 1&,\enspace n\in\mathbb{Z}^{-}\setminus\{-1\}\,.\end{cases}$ (Eq. 3.5.1)
@@
$\lfloor 0\rceil!=\lfloor-1\rceil!=1\,.$ (Eq. 4.1.1)
@@ 4.4 Roman-like definition step 1: $\eta(n)$
In Subsections˜4.4, 4.5, 4.6 and 4.7 we aim to unite the two cases of the original definition of the Roman factorial, expressed as
@@ 4.5 Roman-like definition step 2: $\theta(n)$
For the second step, let’s observe that the factorial in the second case is inverted, while the factorial in the first case remains unchanged. In other words, the factorial of $n$ or (n1)(-n-1) is raised to the power of 1 or 1-1, respectively:
@@
To express this change, we need a function for the exponent that equals 1 when $n$ is non-negative and -1 when $n$ is negative. Another way to describe this requirement is the output pattern [1,1,1][-1,1,1].
@@
In the second case of Eq. 4.5.3, there is a negative sign before the $n$ in the term (n1)(-n-1). This ensures that the factorial calculated is always positive, because if $n<0$, then n>0-n>0. In contrast, in the first case, the $n$ inside the factorial is written without any sign (implying the positive sign).
@@
We are nearly at the point where both cases have identical expressions. The only remaining difference is the 1-1 inside the factorial of the second case, which we’ll address next.
@@ 4.7 Roman-like definition step 4: $\xi^{\prime}(n)$
Let’s investigate the 1-1 that appears only when $n\in\mathbb{Z}^{-}$. We can begin by expanding the first case as shown below:
@@
This marks the completion of one of our objectives. To sum up, we established a collection of foundational functions, which enabled us to express the definition of the Roman factorial concisely and efficiently. In the following section, we will introduce another set of F.F. which will be used alongside those defined in Section˜3 in order to build advanced factorial definitions. For example, in Section˜6, we aim to create a factorial similar to the Roman factorial, but expressed as a non-recursive product.
@@ 5.3 The function $Q(n)$
The function $Q(n)$ represents the signsign function, which exhibits the output pattern [1,0,1][-1,0,1]. It returns the sign of a number, like $\theta(n)$, but it equals 0 if that number is 0. Its piece-wise definition is this:
$Q(n)=\begin{cases}\,1\,,&n>0\\ \,0\,,&n=0\\ -1\,,&n<0\end{cases}\quad,\enspace n\in\mathbb{R}\,.$ (Eq. 5.3.1)
@@
Next, we will attempt to introduce another F.F. in the definition of $Q(n)$. Let’s start by examining the second case: we know that $0=\theta(0)-1$, so we can substitute that expression in the definition. Additionally, let’s rewrite the first case as $\theta(n)-0$, in order to highlight a subtraction from $\theta(n)$:
@@
This function is useful because it behaves identically to $sign(n)$, producing +1+1 or -$1$ for positive or negative inputs respectively, and 0 for the input 0. However, $Q(n)$ is defined using foundational functions, and henceforth, it will be used in place of $sign(n)$.
@@ 5.4 The function $Q^{\prime}(n)$
An easy approach to achieve this goal using the F.F. is to take the absolute value of $Q(n)$. Since $Q(n)$ has the output pattern [1,0,1][-1,0,1], we see that $|Q(n)|$ will have the pattern $[1,0,1]$. That is our desired set of outputs, so we can define $Q^{\prime}(n)$ as the absolute value of $Q(n)$:
@@
Another approach is to subtract $\Theta(n)$ from 1, which is equally valid. Generally, the Boolean-like foundational functions presented in Section˜5 can be combined in many ways to achieve a certain output pattern. Below you will find more alternative definitions of this function:
@@ 5.5 The function $\Psi(n)$
Additionally, we can interpret $\Psi(n)$ as a function that maps 0 to $1$ while leaving all other values of $n$ unchanged. The simplicity of this function makes it useful for various mathematical operations, but it’s most useful in defining the next F.F., $\Phi(n)$.
@@ 5.6 The function $\Phi(n)$
This works well, except for the case when $n=0$: $\Phi(0)=0^{0}$, which is problematic because $0^{0}$ is undefined [^16] The indeterminate form $0^{0}$ does not have a universally agreed-upon value, although it is given the value 1 in a lot of contexts [3]. More about $0^{0}$ in Addendum 9.6.. Fortunately we can solve this problem by using $\Psi(n)$. We end up with this universal definition of $\Phi(n)$:
@@ 5.7 Synopsis
In short, this section is about presenting another set of 5 foundational functions. They were all expressed in terms of other F.F., and along with those in Section˜3 they constitute a set of Boolean-like functions that are instrumental to this paper.
In the following table, we list every foundational function defined in Sections˜3 and 5 for $n\in\mathbb{R}$:
$\delta(n)$ $=\lfloor n\rfloor+0.5$ $[-,+,+]$ (Eq. 3.2.3)
@@
For example, the output pattern of $\xi^{\prime}(n)$ is $[1,0,0]$. That can be read as the binary number $100_{2}$, which is equal to 4 in base 10. Many F.F. exhibit this behavior, but those with outputs like 1-1 or $n$ are excluded.
@@
Table Tbl. 5.7.2: $F.$ outputs in binary
@@ 6.1 Introduction
This section is about generalizing the $\prod$-product definitions of the Roman factorial, in the same manner as it was done in Section˜4.
In Subsection˜2.1, we introduced two formulations of the factorial function: one as a rising product and the other as a falling product. These expressions are reiterated below:
$n!=\prod_{k=1}^{n}k\,=1\cdot 2\cdot 3\cdots(n-2)(n-1)\>\!n\,,\quad n\in\mathbb{Z}^{+},$ (Eq. 2.1.1)
@@
We will follow a methodology akin to that presented in Section˜4: utilizing F.F., we will formulate a piece-wise definition of the factorial. However, at present, we lack a $\prod$-product representation for negative Roman factorials, either as an ascending or descending product.
@@ 6.2 Rising product expression
$n$ -7 -6 -5 -4 -3 -2 -1
$\lfloor n\rceil!$ $\nicefrac{{1}}{{720}}$ -$\nicefrac{{1}}{{120}}$ $\nicefrac{{1}}{{24}}$ -$\nicefrac{{1}}{{6}}$ $\nicefrac{{1}}{{2}}$ -1 1
@@
Indeed, negative Roman factorials seem to be the product of consecutive reciprocals of integers, albeit with an offset. Despite the fractions getting smaller with each increment, since the terms start at 1-1 and end at $n$, we designate this a rising product. Disregarding the offset for now, this pattern can be expressed as follows:
$\lfloor n\rceil!=\prod_{k=1}^{-n}\cfrac{1}{-k}\;,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.2.1)
@@
There are two approaches to address this issue: one is to adjust the upper limit of the product, subtracting 1 from it to multiply by one less term. The other involves multiplying the existing product by $n$, which cancels out the term causing the offset. Both methods work equally well, although the first idea utilizes the empty product [^17] For more information about the empty product, check Addendum 9.2. when n=1n=-1.
@@
$\lfloor n\rceil!=n\cdot\prod_{k=1}^{-n}\cfrac{1}{-k}\>,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.2.2)
@@ 6.3 Falling product expression
From our previous analysis, we observed that the negative Roman factorial of $n$ can be represented by a product consisting of (n1)(n-1) terms. Thus, the upper and lower limits of the product must be set accordingly. A possible definition expressing the aforementioned products is as follows:
$\prod_{k=1}^{-n-1}\cfrac{1}{n+k}=\cfrac{1}{n+1}\cdot\cfrac{1}{n+2}\cdots\cfrac{1}{-2}\cdot\cfrac{1}{-1}\>,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.3.1)
While accurate and applicable to all negative integers, for n=1n=-1, the use of the empty product is necessary again as the upper limit becomes smaller than the lower one. It is preferable to have a definition with the same lower product limit $k=0$ as in Eq. 2.1.2. So, we can multiply by one extra term and subsequently cancel it out, as demonstrated below:
$n\cdot\prod_{k=0}^{-n-1}\cfrac{1}{n+k}=\cfrac{n}{n}\cdot\cfrac{1}{n+1}\cdots\cfrac{1}{-2}\cdot\cfrac{1}{-1}\>,\quad n\in\mathbb{Z}^{-}.$ (Eq. 6.3.2)
Eq. 6.3.2 produces the desired values while also avoiding using the empty product when n=1n=-1. Therefore, we will adopt this definition for the falling product expression:
@@
In summary, in Subsections˜6.2 and 6.3, we’ve established two expressions describing negative Roman factorials as rising or falling products. They are consolidated below:
@@
$n\in\mathbb{Z}^{+}$ $\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Table Tbl. 6.4.1: Roman factorial as a rising or
falling product (not including $n=0$)
@@
$n\in\mathbb{Z}^{+}_{0}$ $\prod_{k=1}^{n}k$ $\enspace\displaystyle\prod_{k=0}^{n-1}(n-k)$
@@
Table Tbl. 6.4.2: Roman factorial as a
rising or falling product
Next, we notice the upper limits of the products. For $n\in\mathbb{Z}^{+}_{0}$, they are either $n$ or n1n-1, and for $n\in\mathbb{Z}^{-}$ they are n-n or n1-n-1.
Essentially, when $n$ becomes negative, a negative sign is applied to ensure that the upper limits remain positive numbers. As seen in Subsection˜4.6, we can replace occurrences of $n$ and n-n with $|n|$. Consequently, the table is modified as follows:
@@
$n\in\mathbb{Z}^{+}_{0}$ $\prod_{k=1}^{|n|}k$ $\enspace\displaystyle\prod_{k=0}^{|n|-1}(n-k)$
@@
Table Tbl. 6.4.3: Roman factorial as a rising or
falling product (generalization: step 1)
Unlike Section˜4, this section generalizes two definitions simultaneously. While it’s possible to separate the procedure for the rising product from the falling one, the steps are identical: combining the two Roman factorial product definitions into a single table allows us to present the generalization process concurrently, unifying the expressions top to bottom.
@@ 6.5 $\prod$-product definition step 2: $\Phi(n)$
In this step, we will introduce the function $\Phi(n)$. As defined in Subsection˜5.6, it takes the form:
@@
Table Tbl. 6.5.1: Roman factorial as a rising or
falling product (generalization: step 2)